ok. i have never understood logarithmic equations
do u mind helping?
4log(1-5x)= 2
i remeber having to do somthing like puting a 4 after log and it cancels and moves to the 2... something like that... can u show me the steps on how to solve logarithms :-]

Found 3 solutions by stanbon, nerdybill, RAY100:
Answer by stanbon(75887)

4log(1-5x)= 2
--------------------
A "log" is an exponent
---
In your problem 4 times the "exponent" = 2
So the "exponent" must be 2/3 or 1/2.
----------------------
So log(1-5x) = 1/2
-----
You might ask: Exponent of what?
Since your problem does not show the base,
the word "log" means the base is 10.
-----------
So 1-5x = 10^(1/2)
----
Solve for "x"
5x = 1-sqrt(10)
x = -0.4325...
===============================
Cheers,
Stan H.

Answer by nerdybill(7384)

4log(1-5x)= 2
log(1-5x)= 2/4
log(1-5x)= 1/2
1-5x = 10^(1/2)
1-5x = 3.1623
-5x = 2.1623
x = -0.432

Answer by RAY100(1637)

remember def of log,,,log(10)a=b if and only if 10^b =a
.
and log a^b =b*log a
.
4 log(1-5x) = 2
.
log (1-5x)^4 =2
.
10^2 =(1-5x)^4
.
100 = (1-5x)^4
.
taking 4th power of both sides,,
.
(100)^(1/4) = 3.162 = (1-5x)
.
3.162 -1 = 2.162 = -5x
.
x= -.4325
.
checking
.
4 log(1-5x) =2
.
4 log (1-5(-.4325) ) =2
.
4log(3.162) =2
.
4*.5 =2
.
2=2,,,,,ok